OpenStudy (anonymous):

f(x) = 4 + 4x^2 − x^4 find inflection points.

OpenStudy (anonymous):

http://www.algebrahelp.com/calculators/equation/ this will give you X

OpenStudy (imstuck):

find the second derivative of this, set it equal to 0 and these will be your points.

OpenStudy (imstuck):

rewrite in descending order to make it a little more digestible!!! lol

OpenStudy (anonymous):

\[12x^2-8\]=0

OpenStudy (anonymous):

\[x=\sqrt{2/3}\]

OpenStudy (imstuck):

The second derivative is\[f''(x)=-12x ^{2}+8\]set it equal to 0:\[0=-12x ^{2}+8\]or better yet:\[0=12x ^{2}-8\]

OpenStudy (anonymous):

then to get y coordinates i plug that into f(x) and I get 6, but its not the right answer

OpenStudy (imstuck):

its not \[\sqrt{\frac{ 2 }{ 3 }}\]it's\[\sqrt{\frac{ 3 }{ 2 }}\]

OpenStudy (anonymous):

why and how do you know which number to bring over first?

OpenStudy (imstuck):

like this:

OpenStudy (imstuck):

wait you're right...I'm wrong! It is your answer, so let's see what's up!

OpenStudy (anonymous):

its 56 over 9 stupid mistake

OpenStudy (imstuck):

\[f''(\sqrt{\frac{ 2 }{ 3 }})=12(\sqrt{\frac{ 2 }{ 3 }})^{2}-8\]\[f''=\frac{ 24 }{ 3 }-8\]That equal 0. Did you get 0?

OpenStudy (imstuck):

wait, what is 56/9?

OpenStudy (imstuck):

I get 0 for the inflection point.

OpenStudy (anonymous):

how

OpenStudy (anonymous):

i gotta leave for class in 8 mins, thanks buddy i'll be on here later asking questions.

OpenStudy (imstuck):

because\[f''(\sqrt{\frac{ 2 }{ 3 }})=12(\sqrt{\frac{ 2 }{ 3 }})^{2}-8\]and\[f''=12(\frac{ 2 }{ 3 })-8\]and\[f''=\frac{ 24 }{ }-8=0\]

OpenStudy (imstuck):

thats over 3. it = 0, test -1 in f'' and 1 in f'' annd you get that the function is increasing to the inflection point

OpenStudy (imstuck):

\[f''=\frac{ 24 }{ 3 }-8\]is\[f''=8-8\]try -1 and +1 in your f'' equation and see what the results are!

OpenStudy (imstuck):

I bet it looks like this:it's increasing on the negative side of 0 and increasing on the positive side of 0

OpenStudy (imstuck):

TY for the medal, even if you are gone from here!